OpenStudy (fanduekisses):

INTEREST HELP!!!: Jacques invests a total of $10,000 in two mutual funds. After 1 yr, one fund produced 12% growth, and the other lost 3%. Find the amount invested in each fund if the total investment grew by $900.

OpenStudy (fanduekisses):

prt + prt = 900 ????

OpenStudy (anonymous):

Theoretically, yes, but we don't know either p.

OpenStudy (anonymous):

You need a second equation. Any ideas?

OpenStudy (fanduekisses):

mhmmm P = I/rt ???

OpenStudy (fanduekisses):

:((

OpenStudy (wolf1728):

prin1 * 1.12 = total 1 prin2 * 1.03 = total 2 prin1 * 1.12 + prin2 * 1.03 = 10,900.00

OpenStudy (fanduekisses):

why 1.12 and 1.03????

OpenStudy (wolf1728):

that's 12% and 3% interest

OpenStudy (anonymous):

Actually it's 1.12 for the 12% growth and 0.97 for the 3% loss. So if we say P1 goes with 12% and P2 goes with -3% then we have \[1.12P _{1}+0.97P _{2} = 10900\]

OpenStudy (anonymous):

But you still need a second equation. Both principles add up to 10000, so \[P _{1}+P _{2}=10000\]

OpenStudy (wolf1728):

Sorry for forgetting the signs :-(

OpenStudy (anonymous):

And you need to solve that system of equations

OpenStudy (fanduekisses):

why 1.12 and not 0.12? I'm so confused right now sorry if it sounds dumb.

OpenStudy (wolf1728):

I'm happy to help you but it seems openstudy keeps bouncing me out of here

OpenStudy (anonymous):

@wolf1728 I'm having that problem as well

OpenStudy (anonymous):

You can do use either and still get the same answer. If you use 1.12 and 0.97, then the taking into account the money already in the account. That's why it adds up to 10900. If you only use 0.12 and -0.03, then you're only accounting for the interest. In that case the equation will be \[0.12P _{1}-0.03P _{2}=900\]

OpenStudy (fanduekisses):

oooooooooh I'm starting to understand it now. ok I'll try solving the system of equations now. :D Thanks.

OpenStudy (wolf1728):

Yeah openstudy ARRGGGHHHH!!!!!! Anyway if we wanted to see how mush 12% interest yields on 10,000 for one year the formula is : principal*(1 + rate )^years 10,000 * 1.12^1 = 11,200

OpenStudy (fanduekisses):

Ohh thank you, so helpful!

OpenStudy (wolf1728):

Yeah, I had a feeling you'd want that 1.12 explained a little bit better. Here's a good page that explains compound interest: http://www.1728.org/compint2.htm

OpenStudy (fanduekisses):

Thank you ^_^